2/3 is what percent of 1/4?
Answer by stanbon(75887)

2/3 is what percent of 1/4?
------
(2/3)/(1/4) = (2/3)(4/1) = 8/3 = 2 2/3 = 2.6666 = 266.66%
-------------
Cheers,
Stan H.
-----------